# line of sight CSI channel state information

### v2

@@ II-B FR3 frequency-dependent features and losses
As discussed in the previous section, empirical field data [5] in InF (dense-scatterer environment) indicate frequency-dependent trends for DS and angular spread (AS). This is not suprising, as frequency-dependence propagation features were measured and revealed in [13] for both outdoors and indoors empirically discovered in what eventually became the FR3 bands of $28$ and $73\operatorname{GHz}$. For modeling the channel path loss over distance, extensive empirical measurements show the close-in free space model with a frequency-weighted path loss exponent (CIF) is more suited for indoors [13], and extends the CI model which has a physics-based close-in free space reference distance as a leverage point for the slope of the exponential path loss, while incorporating the frequency dependence feature of path loss observed indoors. The multipath temporal and spatial statistics at the lower part of the FR3, e.g. $7\operatorname{GHz}$, varies considerably compared to the higher part of FR3, e.g. $16$-$24\operatorname{GHz}$, which confirms less multipath dispersion in both time and space at higher mmWave frequencies [1]. Perhaps most importantly, rain attenuation and foliage losses vary significantly across the FR3. Following ITU-R recommendation (P.838838838838-$3$), the specific attenuation at $7\operatorname{GHz}$ for heavy rain (corresponds to rain rate of $8\operatorname{mm/hr}$) is $0.04\operatorname{dB/km}$, whereas at $24\operatorname{GHz}$, the specific attenuation is $1.16\operatorname{dB/km}$, i.e. a gap of $1.16\operatorname{dB}$ difference over $1\operatorname{km}$. Following Weissberger’s model, the foliage loss difference at $24\operatorname{GHz}$ relative to that at $7\operatorname{GHz}$ is $14.52\operatorname{dB}$ at $100\operatorname{m}$. In particular, the foliage loss is $34.66\operatorname{dB}$ at $7\operatorname{GHz}$ over $100\operatorname{m}$, whereas it is $49.18\operatorname{dB}$ at $24\operatorname{GHz}$ over $100\operatorname{m}$. However, the link can be improved to extend the transmitter’s coverage area via coherent multi-beam combining [12].
III Spectrum Agility & Coverage-Rate Tradeoffs
III-A Spectrum Agility for Coverage-Rate Tradeoffs
Under a constant EIRP with shrinking aperture assumption, the lower FR3 frequencies provide better coverage, as detailed in Section II. If instead the antenna physical area is held fixed, the extra array gain available at mmWave more than compensates for the additional free space pathloss. A fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ? A particularly interesting ray tracing simulation was conducted in [14] to study the coverage and rate statistical behaviour across the FR3 spectrum. In terms of rate, with no blockages, and as expected, it was shown that $24\operatorname{GHz}$ usage dominates in terms of rate for outdoors. Interestingly, when a 3GPP blockage model B was used, it was shown that there was a fallback need towards $12\operatorname{GHz}$ and $6\operatorname{GHz}$ for maximum rate. But even more, and making use of the the comprehensive and detailed penetration models and findings from NYU WIRELESS [1], it is observed that while materials like glass and wood exhibit relatively permeable characteristics at FR3 frequencies, others, e.g. low-emissivity glass, are not permeable. A particularly intriguing observation is that concrete, which is relatively permeable at sub-$6$ $\operatorname{GHz}$ frequencies, becomes completely impermeable at FR3. Additionally, as reported in [14], the indoor user rate distribution indicates that users operating at $24\operatorname{GHz}$ will need to fallback to lower frequencies, which means that, if we really want to leverage the large FR3 spectrum, frequency hopping mechanisms between FR3 bands should be designed to account for the various blockages and penetration losses. The diversity and hopping mechanism naturally pushes forward the need for a spectrum agility in the FR3. Besides, the agility also inherently secures certain transmissions from malicious users and eavesdroppers. A natural question then to ask would be what hopping mechanism should be adopted between the FR3 bands in order to be spectrally efficient and agile?
III-B Incumbents: Coexistence & Opportunity
The primary challenge of utilizing FR3 for $6$G lies in regulatory issues. The spectrum is densely populated with both civilian and government incumbents for various applications beyond fixed and mobile wireless access, like meteorology, radio astronomy, and maritime radio navigation. In contrast to $5$G, one would need to pay special care towards such incumbents. Even if regulators reach consensus on spectrum availability and licensing frameworks, the most technically demanding aspect will be determining how to share this spectrum without interfering with existing users. Note that satellite users are already spectrum-sharing amongst themselves. Vice-versa, we can also consider a fundamental question that is: are commercial services and malicious users able to harm satellites? A possible approach would be through interference nulling to guarantee protection of satellite uplink (UL) communications without sacrificing much the terrestrial downlink (DL), which translates into a greater chance of co-existence. Sharing spectrum efficiently while avoiding interference, especially in cases where users operate close to one another, is a technical challenge that requires effective coordination mechanisms. Relocating incumbent users can be expensive, often eclipsing the auction value of the spectrum itself. Inspired by spectrum refarming mechanisms, dynamic spectrum sharing (DSS) can ensure backward compatibility with older generations in a dynamic way. More specifically, DSS allows service providers to bypass a static cut between different generations in frequency, allowing for wiser spectral utilization. Similarly, DSS can allow a dynamic smart way of sharing spectrum between $6$G users and incumbents within the band, but this would be very challenging.
IV Massive MIMO, Compact Arrays & Power Considerations
Doubling the frequency implies a loss by a factor of $4$ in Watts. However, this loss can be regained by filling the same effective area, covering the space, which also means more compact (but not bigger) arrays, which is referred to as antenna gain. Given half a wavelength spacing, then multiplying the carrier frequency by $N$ necessitates $N^{2}$ more antennas to cover the aperture area. For FR3, transitioning from $3.5\operatorname{GHz}$ to $7\operatorname{GHz}$ faces a $6\operatorname{dB}$ free space path-loss, but also a $6\operatorname{dB}$ antenna gain assuming the same aperture size, enabling similar coverage to lower frequencies with higher efficiency. Meanwhile, the coherence distance shrinks with higher frequencies, which implies that the minimum distance over which signals are uncorrelated in phase and amplitude also decreases, giving rise to the need of packing up more antennas in a given aperture. The challenge remains in the NLoS case, especially for neutral host, where there are much more severe penetration losses at higher frequencies compared to lower counterparts, as mentioned in Sections II and III, in addition to diffraction losses which scale as $10\log f$.
@@ IV Massive MIMO, Compact Arrays & Power Considerations
It is worth noting that new approaches have emerged for the design of ultra-wide band (UWB) antenna arrays whereby mutual coupling (MC) is intentionally introduced between different elements to expand the bandwidth. As far as massive MIMO (mMIMO) is concerned, it was shown that compact mMIMO can open extra degrees-of-freedom in frequency domain [15]. Specifically, MC between array elements can be leveraged in tightly coupled colinear antenna arrays to expand the bandwidth. While interesting to attain higher rates, how can sensing be performed under intentionally induced MC in compact mMIMO designs, especially given its frequency dependence across the UWB?
IV-B Indoor reflection modeling for extended targets
The reflections arising from a single target share a common relation simply because they arise from the same target, i.e. extended target (ET). ETs can reveal information besides location, e.g. size, shape, and orientation, which can enable use cases like gesture recognition. In fact, ET modeling is particularly relevant to indoors owing to its property of rich multipath reflections. Therefore, more modeling efforts should be conducted to characterize ET on lower and higher counterparts of FR3. Even more, as alluded to in Section IV-A, and besides UWB, mMIMO can allow resolving ET paths in spatial domain.
IV-C mMIMO architectures: Fully-digital vs. hybrid
As number of antennas grow large to support various FR3 ISAC applications, which will heavily depend on SE and multi-band operations, a fully-digital scheme may seem complex, due to high power consumption issues. For ISAC, one can consider a fully digital array employing low-resolution digital-to-analog and analog-to-digital converters to trade-off quantization noise with power consumption. Therefore, we can ask what is an optimal fully digital design that can balance sensing tasks with reasonable power consumption ? In contrast, hybrid beamforming seems like a good tradeoff to alleviate the power consumption problem, by trading-off some beamforming and beam selection gains, which, in turn has negative reverberation effects on the received SINR and SE for communications, in addition to detection and localization accuracy for sensing. Therefore, what is the optimal analog beamforming number of sub-panels to strike such trade-offs in hybrid design for sensing and communication (S&C)?
IV-D Coherence time across FR3
Higher frequencies naturally reduce the coherence time, which in turn increases the complexity of channel estimation and makes spatial multiplexing more challenging. An inherent property of FR3 is that the coherence time can significantly vary across FR3, especially in mobile and highly dynamic environments. The coherence block shrinks at higher frequencies, hence the channel outdates faster. In turn, this can limit the duration available for multiplexing data, which can also limit the number of terminals that can be multiplexed simultaneously due to constraints on orthogonal pilot sequences fitting within the coherence time. Within the coherence time, we also know that the achievable rate grows as $\min\left(N_{T},N_{R}\right)\log_{2}(1+\operatorname{SNR})$ per unit of frequency. Hence, mMIMO can allow for more data multiplexing in an environment where less coherence time is available, especially at the higher end of FR3 bands. However, special attention is needed for channel state information acquisition. A question here could be: Should higher FR3 frequencies be dedicated to less mobile environments, while lower FR3 bands to highly dynamic environments ?
IV-E RF frontend efficiency and array integration limits
Energy efficiency of power amplifiers can becoem a challenge in the FR3 band, so transmit side losses can limit the link budget advantages. Particularly, packing large numbers of elements into a fixed aperture soon hits practical barriers due to thermal dissipation, EIRP limits, and fronthaul capacity. Therefore, from array design and RF frontend perspective, how can future FR3 radios achieve high energy efficiency in their power amplifiers while packing dense antenna arrays under real world thermal, EIRP, and processing limits, without giving up the very coverage-rate gains the band is meant to provide? Waste factor can serve as a fundamental metric towards the design.
V Multiband sensing for FR-3 ISAC
ISAC use cases planned for FR3 mean that channel models must support S&C functions. Multi-band sensing for future systems can be realized through non-contiguous aggregation, whereby different frequency blocks are spread across different bands or parts of the FR3 spectrum.
Next, we provide an insightful analysis of some aspects of multiband sensing. As frequency increases, there is typically more spectrum available, and more antennas can be packed per unit of space. Thus, we assume that the bandwidth and antenna gain is proportional to the center frequency. It can be verified that when the frequency is doubled, the signal-to-noise ratio (SNR) and the Cramér-Rao bound (CRB) drop by a factor of $2$. An example is provided in Fig. 2. The drop of SNR at higher frequencies shifts the waterfall region, and the lower frequencies have more margin to accurately start estimating the delay as predicted by the CRB. Conversely, since the CRB of the higher frequency system is smaller, it means that after the waterfall region, the higher frequency achieves a more accurate estimate.
Our analysis reveals two insights for leveraging multi-band sensing. The first is to obtain estimates individually and combine via a maximum ratio combining type approach. Specifically, one can assign more weights to estimates obtained at higher frequencies. The second is relevant for bistatic sensing, where we can rely on higher FR3 frequencies to obtain synchronization via LoS path due to the need of accurate timing, but then we can tolerate to fallback to lower frequencies, which we know can cover more targets. Based on this scheme, a good time difference of arrival location-based method can be developed leveraging the entire multi-band information on FR3.
VI Key FR3 features in $6$G Standards
This section explores key modifications envisioned for next-generation $6$G cellular standardization to unlock the full potential of spectrally agile FR3 radios. Specifically, we examine both 3GPP innovations—focusing on radio frame enhancements tailored for spectral agility—and new O-RAN-based features, particularly in O-RAN open radio unit (O-RU) designs, that facilitate intelligent spectrum utilization.
3GPP
The challenges posed by FR3, e.g. wider bandwidths, non-contiguous carriers, increased propagation losses, and dynamic spectrum environments, necessitate innovations in frame design to ensure low-latency, high-reliability, and efficient multi-band operations. Below, we outline key advancements required to enhance spectral agility, synchronization, and resource management in next-generation wireless systems:
  • Multi-band aggregation, i.e. combining separate, non-contiguous spectrum blocks to optimize the balance between performance and complexity, enhancing both S&C capabilities.
  • Dynamic slot configuration for multi-band coordination, which implements AI-driven adaptive slot structures that dynamically adjust frame configurations across multiple bands. The dynamic strategy ensures optimized latency, throughput, and reliability by tailoring subframe allocations to each band’s propagation characteristics.

### v3

@@ II-D FR3 frequency-dependent features and losses
Frequency-dependent propagation features were measured and revealed in [11] for both outdoors and indoors empirically discovered in what eventually became the FR3 bands of $28$ and $73\operatorname{GHz}$. For modeling the channel path loss over distance, extensive empirical measurements show the close-in free space model with a frequency-weighted path loss exponent (CIF) is more suited for indoors [11], and extends the CI model which has a physics-based close-in free space reference distance as a leverage point for the slope of the exponential path loss, while incorporating the frequency dependence feature of path loss observed indoors. Perhaps most importantly, rain attenuation and foliage losses vary significantly across the FR3. Following ITU-R recommendation (P.838838-33), the specific attenuation at $7\operatorname{GHz}$ for heavy rain (corresponds to rain rate of $8\operatorname{mm/hr}$) is $0.04\operatorname{dB/km}$, whereas at $24\operatorname{GHz}$, the specific attenuation is $1.16\operatorname{dB/km}$, i.e. a gap of $1.16\operatorname{dB}$ difference over $1\operatorname{km}$. Following Weissberger’s model, the foliage loss difference at $24\operatorname{GHz}$ relative to that at $7\operatorname{GHz}$ is $14.52\operatorname{dB}$ at $100\operatorname{m}$. In particular, the foliage loss is $34.66\operatorname{dB}$ at $7\operatorname{GHz}$ over $100\operatorname{m}$, whereas it is $49.18\operatorname{dB}$ at $24\operatorname{GHz}$ over $100\operatorname{m}$. However, the link can be improved to extend the transmitter’s coverage area via coherent multi-beam combining [12].
It is critical to realize that the 3GPP channel modeling efforts in TR 38.901 [6] have up until now ignored the frequency-dependent channel characteristics for most temporal/spatial statistics between 66-$100$ GHz, hence these real-world data are vital for informing a new version of the 3GPP channel model in coming years [4, 5].
Figure 3: Rate-coverage analysis over 77 GHz, $14$ GHz, $18$ GHz and $24$ GHz.
TABLE II: Pros and cons of exploiting the 77-$24$ GHz FR3 spectrum for 66G
Pros / Opportunities Cons / Challenges & Research gaps
Wider bandwidth than FR1 but lower pathloss than FR2. Balances coverage with capacity: outdoor-to-indoor reach that mmWave lacks, but higher throughput than FR1. The centimeter wavelengths enables large antenna, unlocking extreme massive MIMO (mMIMO). Narrower beams than FR1 and richer multipath than FR2 make FR3 attractive for ISAC. Compact high-gain arrays possible on drones. Compact arrays with mutual coupling for tightly coupled colinear antenna arrays can expand bandwidth [13]. Rain/atmospheric loss lower than mmWave. ”Bridge band” for sensing, which helps in mitigating Doppler mismatch between FR1+FR3 or FR3+FR2. High-resolution ISAC: 100100400400 MHz chunks, good for multi-target tracking and imaging. Multi-band capabilities yielding a diverse view of the environment. Coverage–capacity sweet-spot: 10 GHz link budget still supports O2I service without mmWave densification. Mid-band inheritance: more loss than FR1, less available bandwidth than FR2. phase noise (PN) at higher frequencies leads to signal-to-noise ratio (SNR) degradations. Aggregating separated sub-bands increases ADC and algorithmic complexity; frequency-dependent channels need characterization. RF front-end parts (including power amplifiers, analog-to-digital converters, and mixers) covering 772424 GHz are immature. Antenna design for entire FR3 is a challenge. Unified channel models for terrestrial, NTN and ISAC links still ongoing. Spectrum coexistence and aggregation policies demand new multidimensional management schemes. Crowded spectrum: lower FR3 hosts fixed links & DoD users, in addition to X-band for synthetic aperture radar (SAR). Upper FR3 overlaps Ku-band. Sharing needed. Incumbent protection needs accurate sensing of satellite terminals. Open research gaps: unified near/far-field models, extremely large aperture array (ELAA)-aware ISAC theory, learning-based spectrum sharing.
III ITU enhancements via FR3
Peak data rates & bandwidth requirements
Shannon’s formula of capacity tells us that doubling the bandwidth offers higher rates than doubling the power. To keep up with the momentum of different cellular generations, carrier bandwidth should be extended. Indeed, code-division multiple access (CDMA) in 33G used a spectrum block (SB) of $5\operatorname{MHz}$, whereas long-Term Evolution (LTE) uses $20\operatorname{MHz}$. For 55G FR1, the SB is $100\operatorname{MHz}$. Moving forward, 66G must open SBs that are at least $400\operatorname{MHz}$ wide to meet the increasing capacity demands of applications like augmented reality (AR)/ virtual reality (VR). For that, if $3$ carriers in FR3 are supported, one can go up to $1.2\operatorname{GHz}$ of possible non-contiguous bandwidth. This numerology seems very reasonable to keep up with the ITU 66G vision to achieve peak data rates of $50\operatorname{Gbps}$ to $200\operatorname{Gbps}$ depending on the scenario. A $1024$- quadrature amplitude modulation (QAM) constellation offers $10\operatorname{bits/DoF}$. With $12$ antennas providing $12$ spatial degrees-of-freedom (DoF) and $1.2\ \operatorname{GHz}$ overall bandwidth, a peak data rate of $144\ \operatorname{Gbps}$ is achieved. To meet and exceed the IMT-2030 requirement of $200\ \operatorname{Gbps}$, the system can:
  • Increase spatial streams (e.g., $16$ antennas to offer $192\operatorname{Gbps}$)
  • Aggregate more carrier bandwidth additional $400\operatorname{MHz}$ to achieve $192\operatorname{Gbps}$.
  • Combine the above to reach $256\operatorname{Gbps}$, surpassing the IMT-2030 vision of $\boldsymbol{200\operatorname{Gbps}}$.
Location accuracy
The multiband results in Fig. 2 show that scaling carrier frequency and bandwidth by $4f_{\mathrm{ref}}$ (with $f_{\mathrm{ref}}=8\ \operatorname{GHz}$) yields a Cramér-Rao bound (CRB) for delay estimation below $10^{-3}$ at $17$ dB SNR, corresponding to sub $\boldsymbol{10}$ cm location accuracy, fully meeting the IMT-2030 target. When penetration loss at high bands is an issue, hopping to lower frequencies can attain better spectral efficiency (SE).
Coverage-rate tradeoffs & Spectrum agility
In Fig. 3, we run a coverage-rate analysis for four distinct frequencies at FR3. We have allocated more bandwidth to higher frequencies, i.e. $7$ GHz, $14$ GHz,$18$ GHz and $24$ GHz are allocated $100$ MHz, $200$ MHz, $300$ MHz, and $400$ MHz, respectively. Analysis shows that up to $200\operatorname{Mbps}$, lower frequencies dominate, with $7\operatorname{GHz}$ sustaining $200\ \operatorname{Mbps}$ for $30\%$ of the time. Beyond $200\ \operatorname{Mbps}$, the optimal band shifts to $18\ \operatorname{GHz}$ (up to $575\operatorname{Mbps}$), and again at $850\operatorname{Mbps}$, demonstrating the need for spectrum agility in FR3. Dynamically hopping between FR3 frequencies can maintain optimal rate-coverage tradeoffs and deliver IMT-2030 user experienced data rate requirements for $\mathbf{\geq 1\ Gbps}$ peak and high average rates.
@@ III ITU enhancements via FR3
Doubling the carrier frequency, e.g. $3.5$ to $7$ GHz, introduces a $6$ dB free space pathloss and shorter coherence distances, requiring $N^{2}$ more half-wavelength elements to fill the same aperture, but this is counterbalanced by an equal $6$ dB antenna gain, while higher frequency NLoS links still face larger penetration and diffraction losses. mMIMO can aid in achieving the spectrum efficiency specified in IMT-2030, given a preliminary study on the number of antennas required. For IMT-2030 connection density, besides resorting to lower bands for coverage, uplink (UL) densification, where operators can add more transmission and reception points, already supported in 3GPP Rel-16, which are physical base stations within a wireless system that act as nodes for transmitting and receiving signals to improve UL coverage and capacity by attaining spatial diversity at the macro level so that if one path is blocked, an alternative can be used instead.
TABLE III: Comparison among FR1, FR2, and FR3
FR Deployment feasibility Coverage Rate Hardware Implications PLE DS AS
FR1 Outdoors Wide Low Low cost Low High Broad
FR2 Indoors & backhaul Narrow High High complexity High Low Narrow
FR3 Outdoors & indoors Medium Mid Flexible Moderate Moderate Moderate
IV Challenges for commercial deployment
In this section, we highlight some challenges for commercial deployment supporting FR3, which include
Non-contiguous bandwidth & dynamic spectrum environment
Non-contiguous carrier aggregation requires higher resolution ADCs and digital-to-analog converters and more sophisticated baseband algorithms, thereby increasing processing complexity and power consumption. Enabling dynamic spectrum access and coexistence with satellite, radar, or other wireless services operating in FR3 is crucial.
Phase noise (PN) at higher frequencies
PN arises as noise occurring from random and rapid phase fluctuations of a given signal. Hence, PN in both clocks and oscillators can create impact the performance, especially at higher frequencies. In sampling, for a given jitter error, higher frequency signals will suffer from higher phase errors, therefore degrading the clock SNR. In analog, a given amount of time jitter also leads to higher PN at higher frequencies. Over FR3, an SNR degradation caused by PN translates to a loss of $10.7\operatorname{dB}$.
Spectrum coexistence
Lower FR3 bands overlap fixed-link and defense allocations, while upper FR3 intersects Ku-satellite services, necessitating accurate incumbent sensing and multidimensional spectrum-sharing policies.
RF frontend, impairments & linearity
PAs, mixers, and tunable filters capable of broadband operation over FR3 are not yet available at commercial scale. In addition, RF front ends, have to meet stringent adjacent channel leakage power ratios over multiple sub-bands within size constraints of user equipments. Moreover, carrier frequency offset and inphase/quadrature imbalance vary across bands, requiring new estimation and compensation techniques. Furthermore, easily tunable filters to operate smoothly across different FR3 bands is a crucial aspect for multi-band operations in order to leverage the full capability offered by FR3. Also, operating in FR3 necessitates advanced electronic components, such as ADCs and PAs, which must support minimal distortion and maintain linear performance over wider bandwidths.
Additional challenges can be found in Table II.
V Open Questions & Potential Solutions
Question 1: How do recent channel-measurement findings at new FR3 mid-band frequencies affect strategies in 5G-NR (and beyond), and what protocol or metric modifications are being considered within 3GPP to maintain link quality and energy efficiency (EE)?
The newly reported ASs in [4] implies that modifications are required for beam management in the new goldilocks FR3 mid-band spectrum. For sensing applications in ISAC, this also means that precise and accurate beam alignment is needed for target detection, localization and tracking. In addition to channel modeling, the call for new metrics and figures-of-merit, such as the Waste Factor, offer potential for EE optimization in 66G wireless systems.
Question 2: What is an optimal fully digital design that can balance sensing and communication tasks with reasonable power consumption ?
Hybrid beamforming seems like a good tradeoff to alleviate the power consumption problem, by trading-off some beamforming and beam selection gains, which, in turn has negative reverberation effects on the received signal-to-interference-plus-noise ratio (SINR) and SE for communications, in addition to detection and localization accuracy for sensing.
Question 3: How to deal with the difference in coherence time between the lower and upper ends of FR3 ?
The lower end frequencies of FR3 can be allocated to more highly dynamic environments, whereas the higher end to less mobile environments.
Question 4: How to promote SE and agility?
Frequency hopping is interesting in NLoS with severe blockages emanating from building materials, where gains depend on the hopping strategy. However, to study hopping gains, practical antenna gains based on aperture size that fits a BS and UE must be used for assessment, as higher band systems employ higher antenna gains than lower bands. Typical additional link budget gains are about $26$ dB at the BS and $13$ dB at the UE for $28$ GHz, compared to $10$ dB and $6$ dB respectively at $7$ GHz. Using realistic gains, rather than the omni-directional assumptions in TR 38.901, will yield much more reasonable performance curves, often showing only marginal improvements for mid-band compared to high-band in such scenarios.
Question 5: Are commercial services and malicious users able to harm satellites on FR3?
Yes. Terrestrial networks can unintentionally disrupt satellites via co-channel interference or even collision risk from dense constellations. Even more, malicious users go further, deliberately jamming, spoofing, hacking or physically attacking spacecraft, potentially denying service leading to seizing control. For that, mitigation should rely on techniques such as interference nulling and dynamic spectrum sharing (DSS), yet none offers total immunity. Consequently, robust coordination rules and continuous security monitoring are essential to ensure satellite safety. Also, harmonious operation with incumbents can be attained via dynamic spectrum access with real-time interference nulling beams and incumbent sensing capabilities.
Question 6: How does FR3 propagation characteristics, effective bandwidth, and antenna array size in the FR3 band bound the positioning error of received signal strength (RSS)-, angle-, and time-based cellular localization with respect to the IMT-2030 111010 cm target?
RSS
-based systems alone, even with several synchronized BSs may not be enough to satisfy the stringent 111010 cm target set by IMT under typical high pathloss and shadowing. Angle-based positioning with large uniform arrays may achieve the required accuracy in LoS but fails in dense urban multipath, revealing strong sensitivity to AS in this spectrum. Therefore, joint multiband processing for time-based and angle-based positioning algorithms may be needed to attain the 111010 cm target. Part of future work should be able to assess the exact localization accuracy over FR3 under its diverse propagation characteristics.
Question 7: What new propagation traits, never captured in FR1/FR2, can FR3 reveal?
One fundamental feature of operating in the FR3, particularly relevant for sensing, is the phenomenon of diffuse scattering [14], leading to dense multipath components (DMC). Surfaces and objects are seen differently over FR3 depending on the wavelength of the incident wave [14]. Within FR3, and for ISAC applications, a target/object interacting with a channel at the lower end of FR3, will naturally have different diffuse scattering properties than that at the upper end of FR3, even when considering the second-order moments of that channel. In turn, the DMC statistics can offer a distinctive feature of identifying and classifying different targets across FR3. In addition, FR3 will accommodate hybrid near-field and far-field beamforming and algorithms as the Fraunhofer distance varies significantly over the lower/upper FR3 spectrum.
Question 8: What key 3GPP and O-RAN innovations enable spectral agile & spectral intelligence FR3 operation in 6G networks?
  • Multi-band aggregation, which stitches together non-contiguous FR3 blocks so radios can treat scattered spectrum as one wide virtual channel, boosting both sensing and data throughput.
  • Dynamic slot configuration using AI, which can reshape frame structures on the fly across multiple bands to hit the sweet spot between latency, reliability, and capacity.